# Nocturnal Eye–Inspired Liquid–to-Gas Phase Change Soft Actuator with Laser-Induced-Graphene : Enhanced Environmental Light Harvesting and Photothermal Conversion

### v1

@@ Abstract
The mobility of robotic systems is fundamentally constrained by their power sources and wiring requirements. While electrical actuation systems have achieved autonomy through battery power and wireless control, pneumatic actuators remain tethered to air supply sources. Liquid-to-gas phase change actuators utilizing low-boiling-point liquids offer a potential liquid, though they typically require substantial thermal input through heating elements that maintain electrical dependencies. External heat sources, particularly light energy, present an alternative for terrestrial applications. However, silicone-based materials, despite their optical transparency, have high volumetric heat capacity and low thermal conductivity, limiting efficient photothermal energy transfer. Previous attempts to address this through graphene or metallic powder incorporation have compromised material properties, including reduced transparency and altered elastic moduli. Inspired by the tapetum lucidum structure found in nocturnal animals’ eyes for efficient light utilization in low-light conditions, this study proposes a novel anisotropic bilayer soft actuator incorporating Laser-Induced Graphene (LIG) on the inner surface of the light-irradiated silicone layer. This creates an anisotropic structure with enhanced photothermal conversion capabilities while maintaining the advantageous properties of silicone. Comparative analysis demonstrates that the proposed actuator exhibits significantly higher photo-induced bending efficiency compared to conventional silicone-based actuators. The response time (T63) improved by 54%, decreasing from 142 seconds for pure silicone to 65 seconds, with recovery response time showing a 48% improvement. This design maintains silicone’s transparency and flexibility while utilizing LIG, which can be fabricated under ambient conditions, facilitating both manufacturing and diverse applications.
@@ 1 Introduction
Currently, most robotic actuators rely on electrical energy. For example, although pneumatic actuators use air as the driving source, they rely on electrical energy for control and to supply pressed air  (?) (?). Therefore, they cannot be considered completely independent of electrical energy. Throughout history, humans have utilized various forms of energy to enrich their lives. Aiming to use energy sources other than electrical energy to drive soft robots can significantly contribute to diversifying actuation methods in robotics.
Several methods have been proposed on using heating methods, such as laser irradiation  (?) (?), micorwave  (?) , heatsink  (?) and electromagnetic  (?) (?). Kim et al. developed an electrothermal pneumatic actuator using nanofiber mats, achieving a lightweight soft actuator that can operate without external devices by utilizing liquid evaporation  (?). However, this approach faces issues such as high power dependency and slow response times.
Various proposals have been made, starting with actuators like pouch motors ( (?))( (?))( (?))( (?))that utilize liquid-to-gas phase change, including pneumatic actuators driven by ambient temperature ( (?)). However, these actuators operate relatively slowly, with many functioning at frequencies below 0.1 Hz. This is due to significant losses in converting environmental energy into driving energy and issues related to heat transfer and thermal conductivity. Sogabe et al. (?) developed a small soft actuator that operates in environments rich in thermal energy, such as warm water. This innovation improved the actuation time for bending to around one second using only environmental energy. However, this actuator leverages the thermal properties of the surrounding water (specific heat capacity of 4186 $\rm{J/kg\cdot K}$ (?)) and is thus limited to underwater use.
To enable use in more diverse environments, systems capable of operating on land are required. A solar-driven actuator presents a promising approach to solving these issues  (?). Shao et al. proposed an actuator that absorbs environmental light energy through blackbody radiation by mixing graphite into silicone, converting it into thermal energy, and driving it using the liquid-to-gas phase change (methanol) (?). However, this method results in the loss of actuator transparency due to the use of graphite, and adding powders like graphite reduces the Young’s modulus, leading to constraints on the actuator’s shape and functionality.
Therefore, developing a mechanism that efficiently absorbs environmental energy, enabling actuation even with highly transparent silicone actuators, is crucial. We introduce mechanisms inspired by biological systems that efficiently absorb environmental energy into the actuator.
@@
Nocturnal animals need to hunt and escape predators under the faint light of the moon and stars. To efficiently transmit this minimal light to the retinal photoreceptors, they possess a special structure called the tapetum lucidum (Fig. 1A). This structure captures light that passes through the photoreceptors and reflects it back to the retinal layer, stimulating the photoreceptors once again. Inspired by the relationship between the tapetum and the retina, which ‘recovers and reuses’ light that would otherwise escape, our study proposes the introduction of a new light-reusing bilayer structure that enables liquid-to-gas phase change, which is inefficient with silicone layers alone.
@@
In this study, we employed to use Laser Induced Graphene (LIG) as the photothermal conversion element for the ”tapetum” layer. In recent years, Laser-Induced Graphene (LIG) has emerged as a prominent research focus in the field of direct laser writing techniques, where organic materials such as polyimide are transformed into graphene-like structures. LIG is characterized by the formation of a porous three-dimensional graphene network structure, produced through rapid thermal carbonization (pyrolysis) via laser irradiation. Lin et al. proposed a method to directly graphenize polyimide film surfaces using $\rm{CO_{2}}$  laser irradiation (?). Subsequently, Chyan et al. (?) demonstrated the versatility of LIG fabrication on diverse substrates including textiles, paper, and food products through repeated laser exposure, thereby accelerating its application in flexible devices. LIG has garnered significant attention as a photothermal conversion material that requires no electrical connections or external wiring. It exhibits efficient light-to-heat conversion properties under laser or lamp irradiation and demonstrates transferability to elastomeric substrates. For instance, Dallinger et al. developed an actuator by transferring LIG onto PDMS-based hydrogel sheets, where the generated heat induced humidity changes in the sheet, resulting in strain-induced bending (?). These characteristics have positioned LIG as a promising material for applications in bioelectronics and soft actuators (?).
@@
Building upon these advances, we focused on expanding LIG’s applications to liquid-to-gas phase change actuators. We added a LIG layer to conventional silicone soft actuators (?) to enable efficient light-driven actuation. We propose a novel anisotropic bilayer structure where a LIG layer is positioned on the light-incident surface inside highly transparent silicone. This design, inspired by the eyes of nocturnal animals, enables efficient light collection and photothermal conversion. The converted thermal energy is transferred to the internal fluid, causing deformation of the silicone actuator and resulting in controlled bending motion. In this study, we present the fabrication process of the proposed actuator and evaluate its performance characteristics. Furthermore, to demonstrate the practical applicability of our system, we developed a multi-finger robot where multiple units of these novel actuators operate synchronously in response to light stimuli.
Figure 1: Concept and background of the proposed actuator. A: Structure of tapetum lucidum in nocturnal animals’ eyes and its application to the proposed actuator concept utilizing enhanced light absorption efficiency and photothermal conversion as driving force. Inset shows an actual photograph of feline tapetum lucidum (white arrow). B: Schematic illustration of the actuator’s driving mechanism and operational principles. C: Photograph of the fabricated actuator prototype. Scale bar = 10 mm. D: Comparative analysis between conventional soft actuators utilizing black materials (e.g., graphene) for photothermal conversion and the proposed actuator design, highlighting key differences.
@@ 2.1 Concept
Figure 1 illustrates the conceptual design of our proposed actuator. While previous phase-change actuators typically employed homogeneous structures, our novel approach focuses on the layered wall architecture to achieve efficient light absorption. As previously discussed, biological eyes utilize a structure called the tapetum to reflect escaped light, thereby increasing photoreceptor stimulation. Similarly, our design incorporates a Laser-induced graphene (LIG) layer to capture light not absorbed by the conventional silicone elastomer layer, converting it to thermal energy through photothermal conversion to efficiently vaporize and expand the internal low-boiling-point liquid (Fig. 1A).
The actuation mechanism of the proposed device is depicted in Fig. 1B. The soft actuator deforms through the vaporization of an internal low-boiling-point liquid, controlled by an external light source. The actuator is designed with asymmetric silicone wall thicknesses, with the thinner wall on the graphene layer side to induce directional bending toward the LIG layer. The operating temperature was set at 33C, above room temperature (25C), and we selected an inert fluorinated liquid (Opteon SF33, Chemours Company, base compound structure: $\rm{Cis-CF_{3}CH=CHCF_{3}}$) as the working fluid.
Figure 1C shows a photograph of the fabricated actuator. The actuator has an overall length of 20 mm with a flattened elliptical cross-section measuring approximately 5 mm along the major axis and 4 mm along the minor axis. As illustrated in Fig. 1D, our proposed actuator achieves a layered structure using LIG without incorporating thermally conductive materials or photothermal converters directly into the silicone walls. This approach not only facilitates efficient heat transfer but also maintains the transparency of the silicone elastomer, enabling broader applications of the elastomer-based system.
@@ 2.2.1 LIG Transfer Process onto Silicone
The LIG patterns are generated on Kapton tape (3 M Polyimide Film Tape 5413, 3 M Company, Minnesota, USA) adhered to a slide glass. The polyimide base of the Kapton tape enables LIG formation through laser irradiation. We utilize a laser system (xTool F1, Makeblock Co., Ltd., Guangdong, China) with specific parameters (200 mm/s speed, 40% power, unidirectional) for LIG patterning.
The LIG sheet was transferred onto silicone (Ecoflex 00-45 Near Clear, Smooth-On, Inc., Pennsylvania, USA). For the transfer process, the LIG sheet, which remained adhered to the slide glass on one side, was positioned and surrounded by a mold. The degassed uncured silicone was then poured into the mold using a vacuum degassing chamber. The curing process was carried out at 60C for 60 minutes. After curing, the mold was separated from the slide glass and the cured silicone was peeled off the Kapton tape, resulting in the successful transfer of the LIG onto the silicone surface, yielding a silicone substrate with an integrated LIG layer.
Figure 2: Fabrication process of the proposed actuator. In step (i), a laser-processed plate with LIG printing on slide glass is integrated with a mold component designed to form the tubular structure. Subsequently, in step (ii), a metallic rod (2.5 mm in diameter) is inserted and secured at the rod fixation port, followed by the injection of degassed Ecoflex 00-45. After curing at 60C for 60 minutes, step (iii) involves carefully removing the mold and peeling the silicone tube from the plate. At this stage, the LIG layer is positioned on the exterior surface; therefore, in step (iv), the tube is inverted. The fabrication process concludes with step (v), where one end of the open-ended tube is sealed with uncured silicone. The actuator becomes operational upon encapsulation with Opteon SF33 and securing the open end with a silicone-coated clip.
@@ 2.3 LIG-transferred soft actuator fabrication process
Figure 2 illustrates the fabrication process of a soft actuator with the LIG layer positioned adjacent to the liquid interface. The fabrication process consists primarily of five steps, designated as (i)–(v). Initially, a LIG pattern (2.6 mm $\times$ 40 mm) is created using the aforementioned laser parameters. The pattern is then positioned and secured within the mold (detailed in Supplementary Data S1) as shown in (i-a). Subsequently, a metallic rod (2.5 mm in diameter) is inserted into the mold, and silicone is introduced and cured as depicted in (ii). In step (iii), the tubular silicone structure with the integrated LIG layer is carefully removed from the mold. Because the LIG layer is initially present on the external surface, the tube is inverted in step (iv) to position the LIG layer on the interior. In step (v), uncured silicone is introduced into one end of the tube to create a seal, resulting in an actuator with one open terminus. The actuator is completed by introducing 40 µL of a low-boiling-point liquid (Opteon SF33, The Chemours Company, Delaware, USA) into the internal cavity and sealing the open end with a silicone-coated clip.
Additionally, the silicone sheets (10 mm square with transferred LIG patterns) were fabricated using the same process with molds adjusted to achieve a thickness of 1.0 mm. These molds were fabricated using a Formlabs Form 3 printer (Formlabs Inc., Somerville, MA, USA). Clear Resin v4 (Formlabs Inc.) was utilized as the resin material for mold production.
To investigate whether placing the silicone layer on the exterior and the LIG layer on the interior would affect LIG’s photothermal conversion efficiency, we examined the temperature response characteristics. A reflector lamp (RF100 V54 WD, 100 V, 54 W; Panasonic Corporation, Osaka, Japan) was used as the light source, and temperature changes on the silicone sheet were measured using a thermal camera (OPTPI64 ILTO15 T090, Optris GmbH, Berlin, Germany).
@@
Figure 3: Durability assessment of LIG-transferred silicone under various mechanical and chemical stresses. (A) Resistance to repetitive stretching deformation, (B) resistance to abrasive wear, and (C) resistance to liquid exposure. No delamination or degradation of LIG was observed in any of the test conditions. Scale bar = 10 mm.
@@ 2.4 Mechanical and chemical durability of LIG-transferred silicone
Figure 3 illustrates the silicone sheet (10 mm $\times$ 10 mm) with a transferred LIG layer, as well as the condition of the LIG after being subjected to 20 cycles of stretching and relaxation, and twisting, respectively. In both cases, no delamination of the LIG layer was observed.
The resistance of the actuator to the liquid was evaluated by immersing the LIG-transferred sheet into a conical tube filled with Opteon SF33. The actuator was agitated for 60 seconds. After the process, the sheet was examined and no detachment or degradation of the LIG layer was observed, confirming its stability under these conditions (Fig. 4).
@@ 3 Heat transfer analysis
Here, $\sigma=5.67037442\times 10^{-8}~{}\mathrm{W\cdot m^{-2}\cdot K^{-4}}$ is the Stefan-Boltzmann constant; $\theta_{s}$ is the temperature of the silicone wall; $A$ is the heat transfer area; and $\epsilon_{h}$ and $\epsilon_{s}$ are the emissivities of the heat source and the silicone wall, respectively ( (?)). The thermal conductivity of the silicone, $\lambda_{s}$, and its thickness, $x$, are given in Table 1  (?).
The actuator’s response time is sufficiently long. Therefore, the thermal Fourier number becomes large ($F_{o}>>1$) and the temperature distribution within the silicone wall is negligible. The energy balance equation for the silicone wall can then be written as
@@
The parameters used in the simulation are summarized in Table 1 (?). (?)
@@
In this simulation, for the purpose of model simplification, several factors such as Opteon SF33 vaporization and silicone elastomer elongation were not taken into account. The numerical values used for verification were obtained from the referenced literature. Additionally, while the radiative heat transfer equations are presented above, in our actual simulation we simplified the heat transfer calculation by considering the heat source as a light source with constant energy $Q_{h}$. The energy absorbed by each material was then calculated using material-specific absorption coefficients($\alpha_{s}$ and $\alpha_{L}$) which are determined experimentally, where $\alpha_{s}Q_{h}$ represents the energy absorbed by the silicone wall and $\alpha_{L}Q_{h}$ represents the energy absorbed by the LIG layer.
@@
Figure 4: Thermodynamic analysis of temperature changes at liquid-solid interface through thermal imaging and simulation A: Experimental setup for thermal imaging analysis.The measurement was conducted with the distance between the light source and the silicone sheet set to 100 mm. B: Real-time thermal imaging of temperature elevation at the liquid-solid interface comparing silicone sheets with- and without-LIG. The LIG-transferred silicone sheet demonstrated more rapid temperature elevation and higher maximum temperature (334.9 K with-LIG vs. 311.5 K without-LIG). C: Comparison between experimental thermal imaging data (solid lines) and thermodynamic simulation results (dashed lines). Time constant (T63) analysis revealed simulated T63 values of 118.1 sec and 53.9 sec for without- and with-LIG groups, respectively, while experimental measurements yielded T63 values of 110.7 sec and 58.3 sec, respectively.
Figure 5: Overview of the proposed actuator and experimental setup. A: Schematic illustration and dimensions of the proposed actuator. The cross-sectional view reveals a distorted elliptical shape, with slight deformation observed at the adhesion interface with the LIG. B: Schematic diagram of the light irradiation experimental setup. The fundamental experiments were conducted with the actuator positioned at a distance of 50 mm from the light source. A 54 W reflector bulb operating at 100 V served as the light source. C: Measurement method of bending angle: The initial state of the actuator before bending was defined as 0 degrees, and the angle by which the tip of the actuator tilted from this position was recorded as the bending angle.
Figure 6: Temporal analysis of bending actuator dynamics. A: Time-lapse photographs illustrating bending deformation during light irradiation with and without LIG. B: Graph of the relative bending angle ratio during light irradiation (n=5). C: Time-lapse photographs depicting the bending recovery process following light deactivation, and D: Corresponding graphical analysis of the relative bending angle ratio changes (n=5). The relative bending angle ratio is defined as the relative angle when the actuator response reaches a plateau (defined as an increase of less than 1 degree over a 20-second proximity period) or the angle in the frame immediately preceding angle reduction, with this value normalized to 1.

### v2

# Nocturnal Eye-Inspired Liquid-to-Gas Phase Change Soft Actuator with Laser-Induced Graphene: Enhanced Environmental Light Harvesting and Photothermal Conversion

@@ Abstract
Robotic systems’ mobility is fundamentally constrained by their power sources and wiring requirements. While electrical actuation systems have achieved autonomy through battery power and wireless control, pneumatic actuators remain tethered to air supply sources. Liquid-to-gas phase change actuators utilizing low-boiling-point liquids offer a potential liquid, though they typically require substantial thermal input through heating elements that maintain electrical dependencies. External heat sources, particularly light energy, present an alternative for terrestrial applications. However, despite their optical transparency, silicone-based materials have high volumetric heat capacity and low thermal conductivity, limiting efficient photothermal energy transfer. Previous attempts to address this through graphene or metallic powder incorporation have compromised material properties, including reduced transparency and altered elastic moduli. Inspired by the tapetum lucidum structure found in nocturnal animals’ eyes for efficient light utilization in low-light conditions, this study proposes a novel anisotropic bilayer soft actuator incorporating Laser-Induced Graphene (LIG) on the inner surface of the light-irradiated silicone layer. This creates an anisotropic structure with enhanced photothermal conversion capabilities while maintaining the advantageous properties of silicone. Comparative analysis demonstrates that the proposed actuator exhibits significantly higher photo-induced bending efficiency than conventional silicone-based actuators. The response time (T63) improved by 54%, decreasing from 142 seconds for pure silicone to 65 seconds, with recovery response time showing a 48% improvement. This design maintains silicone’s transparency and flexibility while utilizing LIG, which can be fabricated under ambient conditions, facilitating manufacturing and diverse applications.
keywords:
Laser-induced graphene, liquid-to-gas phase change actuator and photothermal conversion
@@ 1 Introduction
Currently, most robotic actuators rely on electrical energy. For example, although pneumatic actuators use air as the driving source, they rely on electrical energy to control and supply pressed air [1][2]. Therefore, they cannot be considered completely independent of electrical energy. Throughout history, humans have utilized various forms of energy to enrich their lives. Aiming to use energy sources other than electrical energy to drive soft robots can significantly contribute to diversifying actuation methods in robotics.
Several methods have been proposed for using heating methods, such as laser irradiation [3][4], microwave [5], heatsink [6] and electromagnetic [7][8]. Kim et al. developed an electrothermal pneumatic actuator using nanofiber mats, achieving a lightweight, soft actuator that can operate without external devices by utilizing liquid evaporation [9]. However, this approach faces issues such as high power dependency and slow response times.
Various proposals have been made, starting with actuators like pouch motors ([10])([11])([12])([13])that utilize liquid-to-gas phase change, including pneumatic actuators driven by ambient temperature ([14]). However, these actuators operate relatively slowly, with many functioning at frequencies below 0.1Hz. This is due to significant losses in converting environmental energy into driving energy, heat transfer, and thermal conductivity issues. Sogabe et al.[15] developed a small soft actuator that operates in environments rich in thermal energy, such as warm water. This innovation improved the actuation time for bending to around one second using only environmental energy. However, this actuator leverages the thermal properties of the surrounding water (specific heat capacity of 4186 $\rm{J/kg\cdot K}$[16]) and is thus limited to underwater use.
For broader environmental applications.., systems capable of operating on land are required. A solar-driven actuator presents a promising approach to solving these issues [17]. Shao et al. proposed an actuator that absorbs environmental light energy through blackbody radiation by mixing graphite into silicone, converting it into thermal energy, and driving it using the liquid-to-gas phase change (methanol)[18]. However, this method results in the loss of actuator transparency due to the use of graphite. Adding powders like graphite reduces Young’s modulus, leading to constraints on the actuator’s shape and functionality.
Therefore, it is crucial to develop a mechanism that efficiently absorbs environmental energy, enabling actuation even with highly transparent silicone actuators. We introduce mechanisms inspired by biological systems that efficiently absorb environmental energy into the actuator.
@@
Nocturnal animals need to hunt and escape predators under the moon’s and stars’ faint light. To efficiently transmit this minimal light to the retinal photoreceptors, they possess a unique structure called the tapetum lucidum (Fig. 1A). This structure captures light that passes through the photoreceptors and reflects it to the retinal layer, again stimulating the photoreceptors. Inspired by the relationship between the tapetum and the retina, which ’recovers and reuses light that would otherwise escape, our study proposes the introduction of a new light-reusing bilayer structure that enables liquid-to-gas phase change, which is inefficient with silicone layers alone.
@@
In this study, we employed Laser-Induced Graphene (LIG) as the photothermal conversion element for the "tapetum" layer. In recent years, Laser-Induced Graphene (LIG) has emerged as a prominent research focus in direct laser writing techniques, where organic materials such as polyimide are transformed into graphene-like structures. LIG is characterized by the formation of a porous three-dimensional graphene network structure, produced through rapid thermal carbonization (pyrolysis) via laser irradiation. Lin et al. proposed a method to directly graphenize polyimide film surfaces using $\rm{CO_{2}}$  laser irradiation[19]. Subsequently, Chyan et al.[20] demonstrated the versatility of LIG fabrication on diverse substrates, including textiles, paper, and food products, through repeated laser exposure, thereby accelerating its application in flexible devices. LIG has garnered significant attention as a photothermal conversion material that requires no electrical connections or external wiring. It exhibits efficient light-to-heat conversion properties under laser or lamp irradiation, demonstrating transferability to elastomeric substrates. For instance, Dallinger et al. developed an actuator by transferring LIG onto PDMS-based hydrogel sheets, where the generated heat-induced humidity changes in the sheet, resulting in strain-induced bending[21]. These characteristics have positioned LIG as a promising material for applications in bioelectronics and soft actuators[22].
@@
Building upon these advances, we focused on expanding LIG’s applications to liquid-to-gas phase change actuators. To enable efficient light-driven actuation, we added a LIG layer to conventional silicone soft actuators[15]. We propose a novel anisotropic bilayer structure with a LIG layer on the light-incident surface inside highly transparent silicone. This design, inspired by the eyes of nocturnal animals, enables efficient light collection and photothermal conversion. The converted thermal energy is transferred to the internal fluid, causing deformation of the silicone actuator and resulting in controlled bending motion. In this study, we present the fabrication process of the proposed actuator and evaluate its performance characteristics. Furthermore, to demonstrate the practical applicability of our system, we developed a multi-finger robot where multiple units of these novel actuators operate synchronously in response to light stimuli.
@@ 2.1 Concept
Figure 1 illustrates the conceptual design of our proposed actuator. While previous phase-change actuators typically employed homogeneous structures, our novel approach focuses on the layered wall architecture for efficient light absorption. As previously discussed, biological eyes utilize the tapetum to reflect escaped light, thereby increasing photoreceptor stimulation. Similarly, our design incorporates a Laser-induced graphene (LIG) layer to capture light not absorbed by the conventional silicone elastomer layer, converting it to thermal energy through photothermal conversion to efficiently vaporize and expand the internal low-boiling-point liquid (Fig. 1A).
The actuation mechanism of the proposed device is depicted in Fig. 1B. The soft actuator deforms by vaporizing an internal low-boiling-point liquid controlled by an external light source. The actuator is designed with asymmetric silicone wall thicknesses, with the thinner wall on the graphene layer side to induce directional bending toward the LIG layer. The operating temperature was set at 33C, above room temperature (25C), and we selected an inert fluorinated liquid (Opteon SF33, Chemours Company, base compound structure: $\rm{Cis-CF_{3}CH=CHCF_{3}}$) as the working fluid.
Figure 1C shows a photograph of the fabricated actuator. The actuator has an overall length of 20 mm with a flattened elliptical cross-section measuring approximately 5 mm along the central axis and 4 mm along the minor axis. As illustrated in Fig. 1D, our proposed actuator achieves a layered structure using LIG without directly incorporating thermally conductive materials or photothermal converters into the silicone walls. This approach facilitates efficient heat transfer and maintains the transparency of the silicone elastomer, enabling broader applications of the elastomer-based system.
@@ 2.2.1 LIG Transfer Process onto Silicone
The LIG patterns are generated on Kapton tape (3 M Polyimide Film Tape 5413, 3 M Company, Minnesota, USA) adhered to a slide glass. The polyimide base of the Kapton tape enables LIG formation through laser irradiation. We utilize a laser system (xTool F1, Makeblock Co., Ltd., Guangdong, China) with specific parameters (200 mm/s speed, 40% power, unidirectional) for LIG patterning[23].
The LIG sheet was transferred onto a silicone (Ecoflex 00-45 Near Clear, Smooth-On, Inc., Pennsylvania, USA). For the transfer process, the LIG sheet, which remained adhered to the slide glass on one side, was positioned and surrounded by a mold. The degassed uncured silicone was then poured into the mold using a vacuum degassing chamber. The curing process was carried out at 60C for 60 minutes. After curing, the mold was separated from the slide glass, and the cured silicone was peeled off the Kapton tape, successfully transferring the LIG onto the silicone surface and yielding a silicone substrate with an integrated LIG layer.
@@ 2.3 LIG-transferred soft actuator fabrication process
Figure 2 illustrates the fabrication process of a soft actuator with the LIG layer positioned adjacent to the liquid interface. The fabrication process consists primarily of five steps: (i)–(v). Initially, a LIG pattern (2.6 mm $\times$ 40 mm) is created using the aforementioned laser parameters. The pattern is then positioned and secured within the mold (detailed in Supplementary Data S1) as shown in (i-a). Subsequently, a metallic rod (2.5 mm in diameter) is inserted into the mold, and silicone is introduced and cured, as depicted in (ii). In step (iii), the tubular silicone structure with the integrated LIG layer is carefully removed from the mold. Because the LIG layer is initially present on the external surface, the tube is inverted in step (iv) to position the LIG layer on the interior. In step (v), uncured silicone is introduced into one tube end to create a seal, resulting in an actuator with one open terminus. The actuator is completed by introducing 40 µL of a low-boiling-point liquid (Opteon SF33, The Chemours Company, Delaware, USA) into the internal cavity and sealing the open end with a silicone-coated clip.
Additionally, the silicone sheets (10 mm square with transferred LIG patterns) were fabricated using the same process with molds adjusted to achieve a thickness of 1.0 mm. These molds were fabricated using a Formlabs Form 3 printer (Formlabs Inc., Somerville, MA, USA). Clear Resin v4 (Formlabs Inc.) was used as the mold production resin material.
We examined the temperature response characteristics to investigate whether placing the silicone layer on the exterior and the LIG layer on the interior would affect LIG’s photothermal conversion efficiency. A reflector lamp (RF100 V54 WD, 100 V, 54 W; Panasonic Corporation, Osaka, Japan) was used as the light source, and temperature changes on the silicone sheet were measured using a thermal camera (OPTPI64 ILTO15 T090, Optris GmbH, Berlin, Germany).
@@ 2.4 Mechanical and chemical durability of LIG-transferred silicone
Figure 3 illustrates the silicone sheet (10 mm $\times$ 10 mm) with a transferred LIG layer, as well as the condition of the LIG after being subjected to 20 cycles of stretching and, relaxation, and twisting, respectively. In both cases, no delamination of the LIG layer was observed.
The resistance of the actuator to the liquid was evaluated by immersing the LIG-transferred sheet into a conical tube filled with Opteon SF33. The actuator was agitated for 60 seconds. After the process, the sheet was examined, and no detachment or degradation of the LIG layer was observed, confirming its stability under these conditions (Fig. 4).
@@ 3 Heat transfer analysis
Here, $\sigma=5.67037442\times 10^{-8}~{}\mathrm{W\cdot m^{-2}\cdot K^{-4}}$ is the Stefan-Boltzmann constant; $\theta_{s}$ is the temperature of the silicone wall; $A$ is the heat transfer area; and $\epsilon_{h}$ and $\epsilon_{s}$ are the emissivities of the heat source and the silicone wall, respectively ([24]). The thermal conductivity of the silicone, $\lambda_{s}$, and its thickness, $x$, are given in Table 1 [25].
The actuator’s response time is sufficiently long. Therefore, the thermal Fourier number becomes large ($F_{o}>>1$), and the temperature distribution within the silicone wall is negligible. The energy balance equation for the silicone wall can then be written as
@@
The parameters used in the simulation are summarized in Table 1[26].[27]
@@
In this simulation, for model simplification, several factors, such as Opteon SF33 vaporization and silicone elastomer elongation, were not taken into account. The numerical values used for verification were obtained from the referenced literature. Additionally, while the radiative heat transfer equations are presented above, in our simulation, we simplified the heat transfer calculation by considering the heat source as a light source with constant energy $Q_{h}$. The energy absorbed by each material was then calculated using material-specific absorption coefficients($\alpha_{s}$ and $\alpha_{L}$), which are determined experimentally, where $\alpha_{s}Q_{h}$ represents the energy absorbed by the silicone wall and $\alpha_{L}Q_{h}$ represents the energy absorbed by the LIG layer.